The perimeter of a rectangle is 48cm. The width is one third the length. Determine the dimensions of the rectangle.
Answer by TimothyLamb(4379)

2L + 2W = 48
W = L/3
3W = L
---
2L + 2W = 48
2(3W) + 2W = 48
8W = 48
8W = 48
---
W = 6
L = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php